Differentiate the trigonometric functions


1. = 2 sec x + tan x



2. y = sin x/ 1 + COS X



3. y = cos x cot x

1
Expert's answer
2022-05-25T09:34:21-0400

1.


"y'=(\sec x+\tan x)'=-\dfrac{1}{\cos^2 x}(-\sin x)+\dfrac{1}{\cos ^2 x}"

"=\dfrac{\sin x+1}{\cos ^2 x}=(\sin x+1)\sec ^2 x"

2.


"y'=(\dfrac{\sin x}{1+\cos x})'=\dfrac{\cos x(1+\cos x)+\sin ^2x}{(1+\cos x)^2}"

"=\dfrac{1+\cos x}{(1+\cos x)^2}=\dfrac{1}{1+\cos x}"

3.


"y'=(\cos x\cot x)'=-\sin x\cot x-\dfrac{\cos x}{\sin^2 x}"

"=-\cos x-\cot x\csc x"